# Towards the classification of maximum scattered linear sets of $\mathrm{PG}(1,q^{5})$

### v2

@@ Abstract
Keywords: Linear set, finite projective line, subgeometry, linearized polynomial, $q$-polynomial
@@ 1 Introduction
The linear set $L_{U}$ is scattered if $\dim_{{{\mathbb{F}}_{q}}}\left(\langle v\rangle_{{\mathbb{F}}_{q^{n}}}\cap U\right)\leq 1$ for any $v\in{\mathbb{F}}_{q^{n}}^{d+1}$ or, equivalently, if it has size $(q^{r}-1)/(q-1)$. A nontrivial scattered ${\mathbb{F}}_{q}$}-linear set $L_{U}$ of $\operatorname{{PG}}(d,q^{n})$ of highest possible rank is a maximum scattered ${\mathbb{F}}_{q}$}-linear set (MSLS for short); in this case $U$ is called a maximum scattered subspace of ${\mathbb{F}}_{q^{n}}^{d+1}$. The rank $r$ of an MSLS depends on $d$ and $n$. It holds $r\leq\frac{(d+1)n}{2}$, see [2], with equality when $(d+1)n$ is even [1, 2, 5]. Two ${\mathbb{F}}_{q}$-linear sets are called equivalent (resp. projectively equivalent) if there exists a collineation (resp. a projectivity) of $\operatorname{{PG}}(d,q^{n})$ that maps one onto the other. A canonical ${\mathbb{F}}_{q}$}-subgeometry of $\operatorname{{PG}}(d,q^{n})$ is a set projectively equivalent to the linear set of rank $d+1$ associated with the ${\mathbb{F}}_{q}$-subspace ${\mathbb{F}}_{q}^{d+1}$ (the set of points with coordinates rational over ${\mathbb{F}}_{q}$}). The relevance of subgeometries lies in the following result.
Theorem 1.1.
In this paper any such pair $(\Gamma,\Sigma)$ is called projecting configuration for the linear set $L_{U}$. The properties of $L_{U}$ can be described in terms of such projecting configuration, the subspace $\Lambda$ being immaterial. The corresponding projection map is denoted by $\wp_{\Gamma}$. In [6], it was proved that for some linear sets the projecting configuration is not unique up to collineations. This issue is further formalized in [4], where the definition of $\operatorname{\Gamma L}$-class of a linear set $L_{U}$ is seen from the perspective of projecting configurations. For a linear set $L_{U}$ of $\operatorname{\Gamma L}$-class one the projecting configuration is unique up to the action of $\operatorname{\Gamma L}(r,q^{n})$; in this case $L_{U}$ is called simple. It has been recently proved [11] that the $\operatorname{\Gamma L}$-class of a MSLS in $\operatorname{{PG}}(1,q^{5})$ not of pseudoregulus type is at most two.
@@
Any linearized polynomial fulfilling (1) is called scattered polynomial. There are only four known families of scattered polynomials which are defined for infinitely many values of $n$ [10, 17, 18, 23, 28]:
@@ 2. $(ii)$
$\operatorname{rk}P=1+\dim\langle P,P^{\sigma},P^{\sigma^{2}},\ldots,P^{\sigma^{n-1}}\rangle$
( [^1] From now on, angle brackets $\langle\ \rangle$ without a field as an index denote the projective closure of a point set in a projective space.). Equivalently, $\operatorname{rk}P$ is the minimum size of a subset $\mathcal{I}\subseteq\Sigma$ such that $P\in\langle\mathcal{I}\rangle$. More precisely, the following result holds.
@@
In [7, 8, 11, 14, 22, 30], several results on the characterization of known maximum scattered linear sets of the projective line have been achieved. These results yield a classification of MSLSs of $\operatorname{{PG}}(1,q^{n})$ for all small values of $n$. In what follows, we collect some of these and outline the current state of the art.
Firstly, since all maximum scattered $\mathbb{F}_{2}$-linear sets of $\operatorname{{PG}}(1,2^{n})$ have size $2^{n}-1$ and are equivalent, they are of pseudoregulus type.
@@
In [7], a characterization result for linear set of pseudoregulus type, obtained via their projecting configurations, is presented.
@@
    the intersection points of $\Gamma$ with $\Gamma^{\sigma^{4}}$, $\Gamma^{\sigma}$, $\Gamma^{\sigma^{3}}$, $\Gamma^{\sigma^{2}}$ are of type AA, $A_{1}$, $B$, $B_{2}$ (cf. Lemma 2.1), respectively;
@@
The investigation is strongly influenced by the rank of the points $A$ and $B$. Thus, in Section 4 the case of $\max\{\operatorname{rk}A,\operatorname{rk}B\}=5$ is studied. Continuing the study started in [22], an algebraic curve $\mathcal{V}$} of degree at most four is associated to $\wp_{\Gamma}(\Sigma)$. Up to some special cases, this $\mathcal{V}$ turns out to be an absolutely irreducible curve of genus at most three. In the general case, the Hasse-Weil Theorem guarantees the existence of a point with coordinates rational over ${\mathbb{F}}_{q}$}, while the remaining cases have been analyzed with a case-by-case analysis. It follows that for $\operatorname{rk}A=5$ or $\operatorname{rk}B=5$ any such MSLS is of LP type.
@@ 3 Characterization by equations
(indices taken mod $5$) a relation containing $v$ and $u_{i}$, $i=0,1,2,3,4$ arises, equivalent to (7). Similarly, to obtain (8) compose $u_{i}=v_{i}-\mu^{q^{is}}v_{i+2}+\lambda^{q^{is}}u_{i+1}$ for $i=0,1,2,3,4$. ∎
@@ 4.1 An algebraic condition for being non-LP type
Case $\operatorname{N}_{q^{5}/q}(\mu)\neq 1$. By Proposition 4.1 ($ii$),
@@ 7 Conclusion
In [22], results were found regarding the maximum scattered linear sets of $\operatorname{{PG}}(1,q^{5})$ not of pseudoregulus type. A necessary and sufficient condition was proven for these sets to be of type LP under the assumption that point $A$, as described in Section 2, has rank 5. In this paper, we proved that such scattered linear sets are always of LP type if the rank of $A$ or $B$ is five. The case in which $\operatorname{rk}A=\operatorname{rk}B=4$ remains open. In this case, we can describe their form and/or the polynomial that defines them. Any linear sets with this shape would be a new type. We have not established their existence and conjecture that they do not exist; in any case, we have ruled out their existence for $q\leq 25$. The following theorem summarizes our work. To provide the reader with more insight, we extend the proof to include a description of the steps that led to the result.

### v3

@@ Abstract
Keywords: finite projective space, linear set, linearized polynomial, scattered polynomial
@@ 1.1 Linear sets in finite projective spaces
The linear set $L_{U}$ is scattered if $\dim_{{{\mathbb{F}}_{q}}}\left(\langle v\rangle_{{\mathbb{F}}_{q^{n}}}\cap U\right)\leq 1$ for any $v\in{\mathbb{F}}_{q^{n}}^{d+1}$ or, equivalently, if it has size $(q^{r}-1)/(q-1)$. A nontrivial scattered ${\mathbb{F}}_{q}$}-linear set $L_{U}$ of $\operatorname{{PG}}(d,q^{n})$ of highest possible rank is a maximum scattered ${\mathbb{F}}_{q}$}-linear set (MSLS for short); in this case $U$ is called a maximum scattered subspace of ${\mathbb{F}}_{q^{n}}^{d+1}$. The rank $r$ of an MSLS depends on $d$ and $n$. It holds $r\leq\frac{(d+1)n}{2}$, see [2], with equality when $(d+1)n$ is even [1, 2, 5]. Two ${\mathbb{F}}_{q}$-linear sets are called equivalent (resp. projectively equivalent) if there exists a collineation (resp. a projectivity) of $\operatorname{{PG}}(d,q^{n})$ that maps one onto the other.
@@
Any linearized polynomial fulfilling (1) is called scattered polynomial. There are only four known families of scattered polynomials which are defined for infinitely many values of $n$ [10, 17, 18, 23, 28]. Of these families, only the following two occur in $\operatorname{{PG}}(1,q^{5})$, the focus of this work:
@@ 2. $(ii)$
1.2 Projecting configurations
A canonical ${\mathbb{F}}_{q}$}-subgeometry of $\operatorname{{PG}}(d,q^{n})$ is a set projectively equivalent to the linear set of rank $d+1$ associated with the ${\mathbb{F}}_{q}$-subspace ${\mathbb{F}}_{q}^{d+1}$ (the set of points with coordinates rational over ${\mathbb{F}}_{q}$}). The relevance of subgeometries lies in the following result.
Theorem 1.1.
In this paper any such pair $(\Gamma,\Sigma)$ is called projecting configuration for the linear set $L_{U}$. The properties of $L_{U}$ can be described in terms of such projecting configuration, the subspace $\Lambda$ being immaterial. The corresponding projection map is denoted by $\wp_{\Gamma}$. In [6], it was proved that for some linear sets the projecting configuration is not unique up to collineations. This issue is further formalized in [4], where the definition of $\operatorname{\Gamma L}$-class of a linear set $L_{U}$ is seen from the perspective of projecting configurations. For a linear set $L_{U}$ of $\operatorname{\Gamma L}$-class one the projecting configuration is unique up to the action of $\operatorname{\Gamma L}(r,q^{n})$; in this case $L_{U}$ is called simple. It has been recently proved in [11] that, for any $n\geq 2$, the $\operatorname{\Gamma L}$-class of an MSLS in $\operatorname{{PG}}(1,q^{n})$ not of pseudoregulus type is at most two.
@@ 1.2 Projecting configurations
$\operatorname{rk}P=1+\dim\langle P,P^{\sigma},P^{\sigma^{2}},\ldots,P^{\sigma^{n-1}}\rangle,$
where, from this point on, the angle brackets $\langle\ \rangle$ without a field as an index will be used to denote the projective closure of a point set in a projective space. Equivalently, $\operatorname{rk}P$ is the minimum size of a subset $\mathcal{I}\subseteq\Sigma$ such that $P\in\langle\mathcal{I}\rangle$. The following result holds.
@@
In [7, 8, 11, 14, 22, 30], several results on the characterization of known maximum scattered linear sets of the projective line have been achieved. These results yield a classification of MSLSs of $\operatorname{{PG}}(1,q^{n})$ for $n\leq 4$. In what follows, we collect some of these and outline the current state of the art.
Firstly, since all maximum scattered $\mathbb{F}_{2}$-linear sets of $\operatorname{{PG}}(1,2^{n})$ have size $2^{n}-1$ and are equivalent, they are of pseudoregulus type, see also [24].
@@
In [7], a characterization result for linear sets of pseudoregulus type, obtained via their projecting configurations, is presented.
@@ 1.3 Outline of the paper
    the intersection points of $\Gamma$ with $\Gamma^{\sigma^{4}}$, $\Gamma^{\sigma}$, $\Gamma^{\sigma^{3}}$, $\Gamma^{\sigma^{2}}$ are of type $A$, $A_{1}$, $B$, $B_{2}$ (cf. Lemma 2.1), respectively;
@@
The investigation is strongly influenced by the rank of the points $A$ and $B$. Thus, in Section 4 the case of $\max\{\operatorname{rk}A,\operatorname{rk}B\}=5$ is studied.
This work builds on the results in [22] concerning maximum scattered linear sets of $\operatorname{{PG}}(1,q^{5})$ that are not of pseudoregulus type. In that paper, a necessary and sufficient condition was proven for these sets to be of type LP under the assumption that the point $A$, as described in Section 2, has rank 5. As in [22], an algebraic curve $\mathcal{V}$} of degree at most four is associated to $\wp_{\Gamma}(\Sigma)$. Up to some special cases, this $\mathcal{V}$ turns out to be an absolutely irreducible curve of genus at most three. In the general case, the Hasse-Weil Theorem guarantees the existence of a point with coordinates rational over ${\mathbb{F}}_{q}$}, while the remaining cases have been analyzed with a case-by-case analysis. It follows that for $\operatorname{rk}A=5$ or $\operatorname{rk}B=5$ any such MSLS is of LP type.
@@ 3 Characterization by equations
(indices taken mod 55) a relation containing $v$ and $u_{i}$, $i=0,1,2,3,4$ arises, equivalent to (7). Similarly, to obtain (8) compose $u_{i}=v_{i}-\mu^{q^{is}}v_{i+2}+\lambda^{q^{is}}u_{i+1}$ for $i=0,1,2,3,4$. ∎
@@ 4.1 An algebraic condition for being non-LP type
Case $\operatorname{N}_{q^{5}/q}(\mu)\neq 1$. By Proposition 4.1 (iiii),
@@ 7 Conclusion
In this paper, starting from the results in [22], we proved that the maximum scattered linear sets of $\operatorname{{PG}}(1,q^{5})$ are always of LP type if the rank of $A$ or $B$ is five. The case in which $\operatorname{rk}A=\operatorname{rk}B=4$ remains open. In this case, we can describe their form and/or the polynomial that defines them. Any linear set with this shape would be a new type. We have not established their existence and conjecture that they do not exist; in any case, we have ruled out their existence for $q\leq 25$. The following theorem summarizes our work. To provide the reader with more insight, we extend the proof to include a description of the steps that led to the result.